# ResSVD: Residual Compensated SVD for Large Language Model Compression

### v2

@@ A.3 More Ablation Results
40% SVD 52326.99 59859.41 47643.04
ASVD [44] 5262.11 8806.33 6522.61
SVD-LLM [41] 13.83 57.07 48.47
ResSVD 13.32 56.25 36.61
ResSVD 12.92 46.93 30.51
50% SVD 130388.72 86721.38 79853.46
ASVD [44] 62726.86 117959.06 77773.84
SVD-LLM [41] 24.05 150.58 141.87
ResSVD 23.06 134.91 87.31
ResSVD 22.41 128.80 83.80
60% SVD 52326.99 59859.41 47643.04
ASVD [44] 16221.43 20119.36 16561.39
SVD-LLM [41] 53.20 378.19 310.17
ResSVD 51.03 357.94 249.36
ResSVD 48.67 323.74 260.06
(b) PLC
@@
40% SVD 52326.99 59859.41 47643.04
ASVD [44] 5262.11 8806.33 6522.61
SVD-LLM [41] 13.83 57.07 48.47
ResSVD 14.33 50.96 46.40
ResSVD 12.92 46.93 30.51
50% SVD 130388.72 86721.38 79853.46
ASVD [44] 62726.86 117959.06 77773.84
SVD-LLM [41] 24.05 150.58 141.87
ResSVD 25.92 147.24 113.76
ResSVD 22.41 128.80 83.80
60% SVD 52326.99 59859.41 47643.04
ASVD [44] 16221.43 20119.36 16561.39
SVD-LLM [41] 53.20 378.19 310.17
ResSVD 60.12 364.38 304.21
ResSVD 48.67 323.74 260.06
A.4 Demonstration of Generated Contents
@@
Table 9 shows the generation contents of models compressed by ResSVD compared to the original model under zero-shot conditions. The results indicate that across various input questions, compressed models consistently produce fluent, coherent, and highly relevant responses. Even at 40%compression ratio, the compressed model can also deliver accurate and relevant answers to common sense questions. These results highlight the effectiveness of ResSVD in achieving substantial compression without compromising performance.
Table 9: Generated contents of LLaMA-7B [38] compressed by ResSVD.
@@
A.5 Computation Complexity
ResSVD decomposes the original weight matrix $\boldsymbol{W}\in\mathbb{R}^{m\times n}$ into two low-rank matrices: $\hat{\boldsymbol{W}}_{u}\in\mathbb{R}^{m\times r}$ and $\hat{\boldsymbol{W}}_{v}\in\mathbb{R}^{r\times n}$. The layer compression ratio $R_{l}$ is computed as $R_{l}=1-\frac{(m+n)r}{mn}$. Under a fixed overall compression ratio $R_{o}$, the relationship between $R_{l}$ and $R_{o}$ is given by $R_{l}=\frac{NR_{o}}{k}$, where $N$ is the total number of layers and $k$ denotes the number of last layers to be compressed.
Given an input activation $\boldsymbol{X}\in\mathbb{R}^{n\times m}$, the original output is computed as $\boldsymbol{Y}=\boldsymbol{W}\boldsymbol{X}$. In the compressed model layer, an intermediate state is first computed as $\boldsymbol{I}=\hat{\boldsymbol{W}}_{v}\boldsymbol{X}$, followed by $\boldsymbol{Y}=\hat{\boldsymbol{W}}_{u}\boldsymbol{I}$. The computational complexity of the original model is $N\cdot\mathcal{O}(m^{2}n)$. For the compressed model, the first $(N-k)$ layers remain uncompressed and retain a complexity of $(N-k)\cdot\mathcal{O}(m^{2}n)$, while the compressed $k$ layers incur a cost of $k\cdot\mathcal{O}(m^{2}r+rnm)$. Thus, the total complexity becomes:
$(N-k)\cdot\mathcal{O}(m^{2}n)+k\cdot\mathcal{O}(m^{2}r+rnm).$ (13)
And the rank $r$ is given by:
$r=\frac{mn(1-R_{l})}{m+n}=\frac{mn(k-NR_{o})}{k(m+n)}.$ (14)
$\boxed{N(1-R_{o})\cdot\mathcal{O}(m^{2}n)}$ (15)
Compared to the original computation complexity $N\cdot\mathcal{O}(m^{2}n)$. This indicates that the total computation cost is reduced proportionally to the overall compression ratio. For example, with $R_{o}=40$%, the compressed model requires only 60% of the original computational cost.
A.6 Limitations
Our strategy is to compress the last few layers of LLMs, which consistently leads to performance improvements. However, for simplicity, we currently apply a uniform compression ratio across all weight matrices within each transformer block. Given that MLP and attention layers may exhibit different sensitivities to compression, assigning module-specific compression ratios in a principled way remains a promising direction for future work.
A.7 Broader Impacts.
By enabling the efficient deployment of LLMs on limited hardware, our method has the potential to broaden access to LLMs. However, we also acknowledge that increased accessibility may lower the barrier to misuse, and thus encourage responsible application in practice.
A.8 Performance on VLMs
We also apply ResSVD to the language part of vision-language models (VLMs) such as LLaVA [26], and evaluate its performance on image captioning (Section A.8.1) and visual question answering (Section A.8.2) tasks. All images used in these evaluations are selected from the ImageNet [8] dataset.
A.8.1 Performance of Image captioning
As shown in Figure 6, LLaVA-1.5-7B compressed with ResSVD under 20% compression ratio is still able to produce image captions that are faithful to the visual content. This indicates that the model retains its ability to interpret and describe images accurately, despite parameter reduction.
Figure 6: Image captioning results of LLaVA-1.5-7B [26] compressed by ResSVD under 20% compression ratio. Captions that accurately describe the image content are highlighted in orange.
A.8.2 Performance of Visual Question Answering
Figure 7 presents the visual question answering results of LLaVA-1.5-7B compressed by ResSVD. The model retains strong performance across these question types, indicating that its multi-modal reasoning capability remains intact despite the compression.
Figure 7: Visual question answering outputs generated by LLaVA-1.5-7B [26] compressed using ResSVD under 20% compression ratio. Questions (Q) and model answers (A) are provided, correct answers are highlighted in orange to emphasize answer quality retention.

### v3

@@ A.3.2 Results on Multiple LLM Families
Original 9.71 15.43 15.52 0.31 0.83 0.68 0.57 0.56 0.77 0.49 0.60
20% SVD-LLM 37.52 40.73 47.25 0.20 0.50 0.55 0.35 0.24 0.64 0.22 0.39
ERC-SVD 35.11 42.77 43.10 0.19 0.54 0.58 0.35 0.25 0.66 0.24 0.40
30% SVD-LLM 72.33 67.68 86.56 0.15 0.42 0.53 0.30 0.19 0.60 0.21 0.34
ERC-SVD 66.89 71.48 82.28 0.16 0.41 0.53 0.30 0.20 0.61 0.22 0.35
40% SVD-LLM 143.28 122.55 162.99 0.12 0.34 0.51 0.28 0.16 0.58 0.22 0.32
ERC-SVD 135.66 113.34 154.32 0.12 0.35 0.52 0.29 0.17 0.57 0.22 0.32
50% SVD-LLM 274.04 226.30 308.11 0.12 0.30 0.50 0.27 0.18 0.55 0.20 0.30
ERC-SVD 247.57 170.82 294.00 0.12 0.30 0.50 0.28 0.18 0.56 0.22 0.31
60% SVD-LLM 462.53 338.39 528.22 0.11 0.27 0.48 0.26 0.18 0.54 0.20 0.29
ERC-SVD 460.73 287.19 522.29 0.11 0.27 0.46 0.27 0.19 0.53 0.20 0.29
A.3.3 Additional Ablation Results
The ablation results for LLaMA-2-7B [touvron2023llama2] under 40%, 50%, and 60% compression ratios are presented in Table 8. A similar trend can also be observed here: incorporating both REC and PLC leads to a substantial reduction in perplexity across all settings.
Table 8: Ablation results of residual compensation and partial-layer compression on LLaMA-2-7B [touvron2023llama2] under 40%, 50%, and 60% compression ratios.
@@ A.3.3 Additional Ablation Results
40% ASVD - - NaN 0.15 0.25 0.50 0.26 0.22 0.52 0.18 0.30
SVD-LLM - - 61.96 0.16 0.35 0.55 0.30 0.20 0.57 0.23 0.34
ERC-SVD 54.19 0.17 0.37 0.54 0.32 0.21 0.57 0.23 0.34
45.13 0.18 0.40 0.55 0.33 0.24 0.61 0.21 0.36
43.19 0.20 0.43 0.57 0.35 0.24 0.63 0.23 0.38
50% ASVD - - NaN 0.13 0.26 0.50 0.25 0.23 0.50 0.20 0.30
SVD-LLM - - 129.71 0.14 0.30 0.50 0.28 0.20 0.54 0.23 0.31
ERC-SVD 126.61 0.14 0.30 0.53 0.29 0.21 0.55 0.24 0.32
117.79 0.13 0.33 0.54 0.29 0.22 0.58 0.22 0.33
100.34 0.14 0.35 0.55 0.31 0.22 0.59 0.22 0.34
60% ASVD - - NaN 0.15 0.25 0.50 0.25 0.23 0.52 0.12 0.29
SVD-LLM - - 263.02 0.14 0.26 0.50 0.27 0.20 0.53 0.21 0.30
ERC-SVD 256.38 0.13 0.26 0.49 0.27 0.20 0.53 0.23 0.30
260.01 0.14 0.29 0.50 0.26 0.18 0.53 0.21 0.30
255.70 0.13 0.29 0.52 0.28 0.21 0.55 0.22 0.31
Table 9: Zero-shot accuracy ($\uparrow$) for seven common sense reasoning datasets on OPT-6.7B, Mistral-7B, and Vicuna-7B under 30% compression ratio.
[…]
OPT-6.7B SVD-LLM 0.27 0.40 0.49 0.33 0.22 0.50 0.20 0.34
ERC-SVD 0.22 0.56 0.61 0.40 0.25 0.69 0.24 0.42
Mistral-7B SVD-LLM 0.13 0.44 0.53 0.30 0.20 0.62 0.20 0.34
ERC-SVD 0.14 0.46 0.58 0.33 0.22 0.63 0.24 0.37
Vicuna-7B SVD-LLM 0.22 0.50 0.56 0.35 0.26 0.63 0.21 0.39
ERC-SVD 0.23 0.51 0.60 0.39 0.30 0.67 0.22 0.42
Table 10: Results on Wikitext-2 under different compression ratios (“Ratio”) of LLaMA-2-7B.
Ratio LLM-Pruner ERC-SVD
20% 8.09 7.63
30% 12.59 10.32
40% 20.36 14.17
50% 40.97 24.26
60% 114.23 58.88
A.3.4 Impact of calibration data
We study the impact of calibration data from two aspects: the number of calibration samples and the choice of calibration datasets. Figure 6 shows the evaluation results under 30% compression ratio. Performance improves slightly with more calibration samples, though the gains remain modest, indicating that ERC-SVD is robust even with limited calibration data. Moreover, Table 11 reports results for different calibration datasets. We also evaluate a mixed dataset, constructed by equally combining the three datasets, to assess its effect on model performance.
Table 11: Perplexity ($\downarrow$) and zero-shot average accuracy ($\uparrow$) of LLaMA-2-7B 20% compressed by ERC-SVD with different calibration datasets.
[…]
WikiText-2 7.63 45.37 14.73 0.48
PTB 9.71 29.43 13.29 0.46
C4 9.44 37.67 11.49 0.47
Mix 10.13 36.53 13.13 0.45
Table 12: Perplexity ($\downarrow$) of LLaMA-2-7B compressed by ERC-SVD and SVD-LLM, followed by quantization with GPTQ-8bit [frantar2022gptq]. Blue arrows within parentheses highlight the relative improvement.
[…]
- Original GPTQ-8bit 5.47 26.79 7.28
30% SVD-LLM GPTQ-8bit 10.67 336.30 34.93
ERC-SVD GPTQ-8bit 10.34 ($\downarrow$3%) 79.05 ($\downarrow$76%) 25.02 ($\downarrow$28%)
Figure 5: Throughput of LLaMA-7B and its 40% compressed versions. The sequence length is 32.
Figure 6: Impact of the number of calibration data samples on LLaMA-2-7B under 30% compression ratio. (Left) Perplexity ($\downarrow$). (Right) Average accuracy ($\uparrow$).
A.4 Comparison with Smaller-Scale LLM
To further assess the effectiveness of ERC-SVD under relatively high compression ratios, we conduct a comparative evaluation of 50% and 60% compressed LLaMA-2-7B against Gemma-2-2B [gemma_2024]. Table 13 presents the performance comparison between the compressed LLaMA-2-7B models and Gemma-2-2B. Under 50% compression ratio (approximately 3.5B parameters), LLaMA-2-7B consistently outperforms Gemma-2-2B across most benchmarks. Even under a more aggressive 60% compression ratio (around 2.8B parameters), the compressed model maintains competitive performance, remaining comparable to Gemma-2-2B.
Table 13: Performance comparison of Gemma-2-2B and LLaMA-2-7B compressed by ERC-SVD under 50% and 60% compression ratios (“Ratio”).
[…]
Gemma-2-2B - 3B 67.78 374.60 70.78 0.13 0.39 0.51 0.31 0.19 0.60 0.23 0.33
LLaMA-2-7B 50% 3.5B 24.26 286.24 100.34 0.14 0.35 0.55 0.31 0.22 0.59 0.22 0.34
60% 2.8B 68.59 991.48 255.70 0.13 0.29 0.52 0.28 0.21 0.55 0.22 0.31
Table 14: Performance (measured by accuracy ($\uparrow$)) of original LLaVA-1.5-7B, and its 20% compressed versions by SVD-LLM and ERC-SVD on VLM benchmarks. The best results are marked in bold.
[…]
- Original 88.2 87.3 85.1 58.17 70.15
20% SVD-LLM 82.5 83.2 77.8 30.68 49.54
ERC-SVD 90.2 ($\uparrow$9%) 87.7 ($\uparrow$5%) 83.1 ($\uparrow$7%) 50.86 ($\uparrow$66%) 69.54 ($\uparrow$40%)
A.5 Comparison with Pruning
Table 10 shows the performance of LLaMA-2-7B compressed by LLM-Pruner [ma2023llm] and ERC-SVD under different compression ratios on the WikiText-2 dataset. It can be observed that ERC-SVD consistently outperforms the pruning method, achieving a perplexity of 58.88 under 60% compression ratio, compared to 114.23 for LLM-Pruner.
A.6 Computation Complexity
ERC-SVD decomposes the original weight matrix $\boldsymbol{W}\in\mathbb{R}^{m\times n}$ into two low-rank matrices: $\hat{\boldsymbol{U}}_{r}\in\mathbb{R}^{m\times r}$ and $\hat{\boldsymbol{V}}_{r}\in\mathbb{R}^{r\times n}$. The layer compression ratio $R_{l}$ is computed as $R_{l}=1-\frac{(m+n)r}{mn}$. Under a fixed overall compression ratio $R_{o}$, the relationship between $R_{l}$ and $R_{o}$ is given by $R_{l}=\frac{NR_{o}}{k}$, where $N$ is the total number of layers and $k$ denotes the number of last layers to be compressed.
Given an input activation $\boldsymbol{X}\in\mathbb{R}^{n\times m}$, the original output is computed as $\boldsymbol{Y}=\boldsymbol{W}\boldsymbol{X}$. In the compressed model layer, an intermediate state is first computed as $\boldsymbol{I}=\hat{\boldsymbol{V}}_{r}\boldsymbol{X}$, followed by $\boldsymbol{Y}=\hat{\boldsymbol{U}}_{r}\boldsymbol{I}$. The computational complexity of the original model is $N\cdot\mathcal{O}(m^{2}n)$. For the compressed model, the first (Nk)(N-k) layers remain uncompressed and retain a complexity of $(N-k)\cdot\mathcal{O}(m^{2}n)$, while the compressed $k$ layers incur a cost of $k\cdot\mathcal{O}(m^{2}r+rnm)$. Thus, the total complexity becomes:
$(N-k)\cdot\mathcal{O}(m^{2}n)+k\cdot\mathcal{O}(m^{2}r+rnm).$ (15)
And the rank $r$ is given by:
$r=\frac{mn(1-R_{l})}{m+n}=\frac{mn(k-NR_{o})}{k(m+n)}.$ (16)
$\boxed{N(1-R_{o})\cdot\mathcal{O}(m^{2}n)}$ (17)
Compared to the original computation complexity $N\cdot\mathcal{O}(m^{2}n)$. This indicates that the total computation cost is reduced proportionally to the overall compression ratio. For example, if $R_{o}=40$%, the compressed model requires only 60% of the original computational cost.
A.7 Demonstration of Generated Contents
@@
Table 15 shows the generation contents of models compressed by ERC-SVD compared to the original model under zero-shot conditions. The results indicate that across various input questions, compressed models consistently produce fluent, coherent, and highly relevant responses. Even at 40% compression ratio, the compressed model can also deliver accurate and relevant answers to common sense questions. These results highlight the effectiveness of ERC-SVD.
Table 15: Generated contents of LLaMA-7B [touvron2023llama] compressed by ERC-SVD.
@@
A.8 ERC-SVD on Vision-Language Model
In this section, we apply ERC-SVD to the language part of vision-language models (VLMs) such as LLaVA [liu2024improved], and evaluate both quantitative and qualitative performance.
A.8.1 Quantitative Results
We compare the performance of LLaVA-1.5-7B [liu2024improved] compressed using SVD-LLM [wang2024svd] and ERC-SVD under 20% compression ratio. We report results on several benchmarks: POPE [li2023evaluating], TextVQA [singh2019towards], and ScienceQA [lu2022learn]. The results, shown in Table 14, indicate that ERC-SVD consistently outperforms SVD-LLM across all benchmarks. Notably, it achieves substantial relative improvements, 66% on TextVQA and 40% on ScienceQA. Moreover, on the POPE-random and POPE-popular subsets, the model compressed by ERC-SVD even surpasses the original LLaVA-1.5-7B.
A.8.2 Qualitative Results
As shown in Figure 7, LLaVA-1.5-7B compressed with ERC-SVD under 20% compression ratio is still able to produce image captions that are faithful to the visual content. This indicates that the model retains its ability to interpret and describe images accurately, despite parameter reduction.
Figure 7: Image captioning results of LLaVA-1.5-7B compressed by ERC-SVD under 20% compression ratio. Captions that accurately describe the image content are highlighted in orange.
Figure 8 presents the visual question answering results of LLaVA-1.5-7B compressed by ERC-SVD. The model retains strong performance across these question types, indicating that its multi-modal reasoning capability remains intact despite the compression.
Figure 8: Visual question answering outputs generated by LLaVA-1.5-7B compressed using ERC-SVD under 20% compression ratio. Questions (Q) and model answers (A) are provided, correct answers are highlighted in orange to emphasize answer quality retention.